# Tree Embedding in High Dimensions:

### v1

@@
Abstract
@@ 1.1 Our Results
Our main contribution is a new tree embedding framework in general metrics that works with wide families of metric decompositions (which we discuss in more detail in Section˜1.2). This framework combining with geometric hashing techniques yields new results for tree embedding in dynamic and MPC settings, which in turn yields rich applications. Our approach may be of independent interest to tree embedding beyond Euclidean spaces. Below, we start with introducing the dynamic result and its applications.
@@
Theorem 1.1 (Dynamic Tree Embedding; see Theorem˜4.2).
@@
For any constant $\epsilon$, our result achieves logarithmic distortion, which is asymptotically optimal [AKPW95, Bar98]. [^3] Our result also works if $\epsilon$ depends on $n$, for example $\epsilon=1/\log n$ (where the $\tilde{O}(n^{\epsilon})$ time term becomes $\tilde{O}(1)$). However, in this parameter regime, our result is not competitive and it is open to find tight tradeoffs. In contrast, two implicit approaches for dynamic tree embedding in $\mathbb{R}^{d}$ yield super-logarithmic distortion. (i) Indyk’s [Ind04] tree embedding result achieves $O(d\log n)$ distortion and can handle point updates in $\textrm{poly}(\log n)$ update time. (ii) The work of Forster, Goranci, and Henzinger [FGH21] maintains a tree embedding of dynamic graphs with distortion $O(\log^{3t-2}n)$ and update time $O(m^{1/t+o(1)}\log^{4t-2}n)$, for $t\geq 2$. The high-dimensional spanner construction due to Har-Peled, Indyk and Sidiropulus [HIS13] gives an $O(t)$-spanner with $\tilde{O}(n^{1+1/t^{2}})$ edges, which can be extended to support point updates in $\tilde{O}(n^{1/t^{2}})$ update, for $t\geq 1$. Combining these two algorithms, i.e., running the graph-based dynamic tree embedding result on the dynamic spanner, leads to sub-optimal distortion guarantees.
@@
In a similar vein, our dynamic tree embedding from  Theorem˜1.1 implies new dynamic algorithms for $k$-median clustering, Euclidean bipartite matching (also known as Earth Mover’s Distance), general matching, and geometric transport. For the sake of brevity, we will only review the first two applications and refer the reader to  Section˜5 for an extended discussion of all applications.
In the Euclidean $k$-median clustering problem, the objective is to select a set of $k$ centers for a given set of $n$ points in $\mathbb{R}^{d}$ so as to minimize the total distance from each point to its nearest center. In the fully dynamic setting, the challenge is to maintain a good approximation (which is a set of $k$ centers) to the optimal clustering while efficiently supporting both insertions and deletions of points. Leveraging the structure of the algorithm from [CALNF+21] within our dynamic tree embedding Theorem˜1.1, we obtain the following guarantees for fully dynamic $k$-median clustering.
Corollary 1.2 (Dynamic $k$-median Clustering; see Corollary˜5.1).
@@
Recently, considerable effort has been devoted to understanding the optimal trade-offs for $k$-median clustering in the dynamic setting [HK20, FLNFS21, lTHS24, BCG+24, BCF25]. Bhattacharya, Costa, and Farokhnejad [BCF25] achieve state-of-the-art guarantees, obtaining an $O(1)$-approximation with $\tilde{O}(k)$ update time. However, this $\tilde{O}(k)$ is not sublinear time in the worst case (i.e., $k=O(n)$). In contrast, our algorithm always runs in sublinear small polynomial update time, at the expense of incurring a logarithmic factor in the approximation ratio. Previously, no known algorithm can achieve such sublinear update time with any $\operatorname{poly}\log n$ ratio.
We further consider the classic Euclidean bipartite matching problem in the dynamic setting. Although this problem has been recently studied in low dimensions [GKP+25], the update time of their approach exhibits an inherent exponential dependency on the dimension. We use Theorem˜1.1 to give the first high-dimensional algorithm for this problem.
Corollary 1.3 (Dynamic Euclidean Bipartite Matching; see Corollary˜5.5).
@@
Theorem 1.4 (MPC Tree Embedding; see Theorem˜4.3).
@@ 1.2 Technical Contributions
We provide a high-level picture of our technical contributions as follows. Then in Sections˜1.2.1 and 1.2.2 we give an overview of proofs, where we talk about the concrete technical challenges and our solution.
@@
To utilize the locality guarantee of Item˜1, we employ a metric decomposition called sparse partitions [JLN+05, Fil24]. Specifically, in Euclidean $\mathbb{R}^{d}$, it serves as a $\tau$-BD such that any radius-$r/\Gamma$ metric ball intersects at most $n^{1/\operatorname{poly}(\Gamma)}$ parts. In our applications, we need to use consistent hashing [CJK+22, FJL+25] which is data-oblivious and time/space efficient construction of such sparse partition. This enables us to precisely simulate our new CKR algorithm in the MPC setting. We further show that the embedding undergoes $\tilde{O}(1)$ expected well structured changes under point insertions and deletions. In turn, our new tree embedding results open a door to rich applications. Notably, we devise new dynamic algorithms based on this new tree embedding maintaining optimal solutions with respect to the dynamic embedding tree.
@@ 1.2.1 Proof Overview: Tree Embedding via General Metric Decomposition
In our modified CKR, apart from the input dataset $P\subseteq V$ and scale $w>0$, our algorithm additionally takes a parameter $\Gamma>0$ as input. As mentioned, this parameter balances the accuracy with the locality. Our algorithm start with imposing an $O(w)$-BD/hashing. Then a natural change we make to original CKR, is to replace finding the minimum $\pi$ value in $B^{P}(p,r)$, with that in the intersecting buckets $\tilde{B}(p,r):=\operatorname{buks}^{P}(B(p,r))$, i.e., $\ell_{p}:=\pi_{\min}(\tilde{B}^{P}(p,r))$, where $\pi_{\min}(S):=\min_{x\in S}\pi(x)$. The other related change is that $r$ is picked uniformly at random from $w/\Gamma/4,w/\Gamma/2]$, scaled by $\Gamma$. Finally, there is also a subtle change to defining tree embedding: for each distance scale $2^{m-i}$, we run the modified CKR not only using the scale $w_{i}=2^{m-i}$, but also a new instance of $O(w_{i})$-BD (over $i$) which may not have any relation to other $2^{j}$-BD’s for $j\neq i$, especially that they are not nested/hierarchical. This means we must also define $\tilde{B}^{P}$ as $\tilde{B}^{P}_{i}$ for scale $w_{i}$, instead of using a universal BD. See Algorithm˜1 for the full description of this process.
@@
To adapt this FRT’s analysis in our case, a major technical challenge is caused by using $\tilde{B}_{i}(\cdot,\cdot)$ to define the labels for each scale $w_{i}$ (instead of using the metric ball $B(\cdot,\cdot)$). This approximate metric ball $\tilde{B}_{i}(p,r)$ is structurally sophisticated and does not share nice geometric properties as in an exact metric ball, especially considering we are working with an arbitrary bounded decomposition. To see this, given that a point $q$ belongs to $\tilde{B}_{i}(p,r)$, it does not imply a closer point $q^{\prime}$ (to $p$) is also in $\tilde{B}_{i}(p,r)$, and this breaks the probability $\Pr[\mathcal{E}_{i}]\leq 1/i$ as in (1). We emphasize that these challenges are independent to the new parameter $\Gamma$, and the technical challenges remain even if $\Gamma=1$.
@@
To resolve this issue, our main insight is that it is possible to define a set of representative points $\operatorname{rep}(p,r)\subseteq V$ for $\tilde{B}(p,r)$ (and $\operatorname{rep}(p,r)$ may not be a subset of dataset $P$), defined such that the algorithm may be roughly interpreted as running on $\operatorname{rep}(p,r)$, and that it behaves similarly enough to the metric ball $B(p,r)$. This holds for any metric decomposition. In particular, it satisfies the following (see Definition˜3.6) and the existence of such representatives is shown in Lemma˜3.7, which is our major technical insight.
@@
In this way, we can use $\operatorname{rep}_{i}^{P}(p,r)$ to replace $B^{P}(p,r)$ when computing the label $\ell^{(i)}_{p}$; specifically, the label $\ell^{(i)}_{p}$ can be equivalently found only in $\operatorname{rep}_{i}(p,r)$, instead of the whole $B(p,r)$, i.e., there is $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ such that $\ell^{(i)}_{p}=\pi_{\min}(\operatorname{buks}^{P}_{i}(\operatorname{rep}^{*}_{i}(p,r)))$ (see Definition˜3.8). Moreover, we can show an analogy to (1): for some scale $w_{i}$, list $\operatorname{rep}_{i}(p,r)$ in non-decreasing order of distance to $p$ as $p_{j}$’s,
$\forall j,\quad\Pr[\pi(p_{j})=\ell^{(i)}_{p}]\leq H_{m_{j}}-H_{m_{j-1}},$ (2)
where $H_{t}:=\sum_{s=1}^{t}1/s$ is the harmonic sum, and $m_{t}:=|\bigcup_{s\leq t}\operatorname{buks}^{P}(p_{s})|$ is the size of the first $t$ buckets (see (17) in Section˜3.2).
In the existence proof of $\operatorname{rep}(\cdot,\cdot)$ (in Lemma˜3.7), we actually give an explicit definition of it (used only in the analysis). This definition is technical, and here we discuss the intuition. A natural attempt to define $\operatorname{rep}(\cdot,\cdot)$ such that it satisfies all three of distinct, monotone and ball-preserving properties, is to let $\operatorname{rep}_{i}(r,p):=B(p,r)\cap\operatorname{buks}_{i}(P)$, which is precisely the first half of the “ball-preserving” bullet. The issue is that it does not satisfy the distinctness (bullet 1). Our definition fixes this by only selecting a subset of $B(p,r)\cap\operatorname{buks}_{i}(P)$, roughly the nearest neighbors from $p$ in each bucket of $B(p,r)\cap\operatorname{buks}_{i}(P)$. This “nearest” property ensures the distinctness (bullet 1). Moreover, it is important to select the nearest neighbor and it does not work if e.g., selecting a point in the buckets with smallest ID. The reason is that nearest neighbor has additional geometric properties that ensure bullet 2 and 3 hold.
@@
Our final analysis deviates from that of FRT, and we indeed need a sum over $i$’s. To avoid the additional $O(\log n)$ factor, we derive a refined upper bound with respect to the scale $w_{i}$, instead of a universal one. The more difficult case is when $w_{i}$ is much larger than $\operatorname{dist}(p,q)$, since if $\ell^{(i)}_{p}\neq\ell^{(i)}_{q}$ happens it incurs a large $w_{i}$ as error, which can be sensitive. For this case, we have an improved bound $O(\operatorname{dist}(p,q))\cdot(H_{\tilde{B}_{i}(p,w_{i})}-H_{\tilde{B}_{i}(p,w_{i}/4)})$ (see Lemma˜3.3).
A natural next step is to cancel the terms $H$ out via a telescoping sum when summing over $i$. However, this cancellation requires us to relate $H_{\tilde{B}_{i}}$ with $H_{\tilde{B}_{j}}$ for $j\neq i$. Although the decompositions in different scales may not be related in any way, we can still use the diameter bound and triangle inequality to show that, when the scales differ by a $\operatorname{poly}(\Gamma)$ factor, the smaller-scale approximate ball $\tilde{B}^{P}$ would be a subset of the larger one, regardless of how the decomposition behaves; namely, $\tilde{B}^{P}_{i+s}(p,w_{i}/2^{s})\subseteq\tilde{B}^{P}_{j}(p,r_{i})$ for any $i,j$ and $s=O(\log\Gamma)$ (see Lemma˜3.4). This helps us to only incur an additional factor of $O(\Gamma\log\Gamma)$ instead of $O(\log n)$.

### v2

@@
Tree Embedding in High Dimensions:
Dynamic and Massively Parallel
Gramoz Goranci Faculty of Computer Science, University of Vienna, Austria (gramoz.goranci@univie.ac.at).    Shaofeng H.-C. Jiang School of Computer Science, Peking University, China (shaofeng.jiang@pku.edu.cn).    Peter Kiss Faculty of Computer Science, University of Vienna, Austria (peter.kiss@univie.ac.at). This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/ESP6088024.    Qihao Kong School of Computer Science, Peking University, China (2300012972@stu.pku.edu.cn).    Yi Qian School of Computer Science, Peking University, China (qianyi@stu.pku.edu.cn).    Eva Szilagyi Faculty of Computer Science, UniVie Doctoral School Computer Science DoCS, University of Vienna, Austria (eva.szilagyi@univie.ac.at).
@@ 1.1 Our Results
Our main contribution is a new tree embedding framework in general metrics that works with wide families of metric decompositions (which we discuss in more detail in Section 1.2). This framework combining with geometric hashing techniques yields new results for tree embedding in dynamic and MPC settings, which in turn yields rich applications. Our approach may be of independent interest to tree embedding beyond Euclidean spaces. Below, we start with introducing the dynamic result and its applications.
@@
Theorem 1.1 (Dynamic Tree Embedding; see Theorem 4.2).
@@
For any constant $\epsilon$, our result achieves logarithmic distortion, which is asymptotically optimal [AKPW95, Bar98]. [^3] Our result also works if $\epsilon$ depends on $n$, for example $\epsilon=1/\log n$ (where the $\tilde{O}(n^{\epsilon})$ time term becomes $\tilde{O}(1)$). However, in this parameter regime, our result is not competitive and it is open to find tight tradeoffs. In contrast, two implicit approaches for dynamic tree embedding in $\mathbb{R}^{d}$ yield super-logarithmic distortion. (i) Indyk’s [Ind04] tree embedding result achieves $O(d\log n)$ distortion and can handle point updates in $\textrm{poly}(\log n)$ update time. (ii) The work of Forster, Goranci, and Henzinger [FGH21] maintains a tree embedding of dynamic graphs with distortion $O(\log^{3t-2}n)$ and update time $O(m^{1/t+o(1)}\log^{4t-2}n)$, for $t\geq 2$. The high-dimensional spanner construction due to Har-Peled, Indyk and Sidiropulus [HIS13] gives an $O(t)$-spanner with $\tilde{O}(n^{1+1/t^{2}})$ edges, which can be extended to support point updates in $\tilde{O}(n^{1/t^{2}})$ update, for $t\geq 1$. Combining these two algorithms, i.e., running the graph-based dynamic tree embedding result on the dynamic spanner, leads to sub-optimal distortion guarantees. Finally, we note that Jayaram, Waingarten and Zhang [JWZ24] gave a close-to-optimal $\tilde{O}(\log n)$-distortion in $\operatorname{poly}\log(n)$ update time, but the maintained tree is not 2-HST (see Section 2 for a formal definition) which is structurally weaker and may not lead to the applications as we obtain in this paper.
@@
In a similar vein, our dynamic tree embedding from  Theorem 1.1 implies new dynamic algorithms for $k$-median clustering, Euclidean bipartite matching (also known as Earth Mover’s Distance), general matching, and geometric transport. For the sake of brevity, we will only review the first two applications and refer the reader to  Section 5 for an extended discussion of all applications.
In the Euclidean $k$-median clustering problem, the objective is to select a set of $k$ centers for a given set of $n$ points in $\mathbb{R}^{d}$ so as to minimize the total distance from each point to its nearest center. In the fully dynamic setting, the challenge is to maintain a good approximation (which is a set of $k$ centers) to the optimal clustering while efficiently supporting both insertions and deletions of points. Leveraging the structure of the algorithm from [CALNF+21] within our dynamic tree embedding Theorem 1.1, we obtain the following guarantees for fully dynamic $k$-median clustering.
Corollary 1.2 (Dynamic $k$-median Clustering; see Corollary 5.1).
@@
Recently, considerable effort has been devoted to understanding the optimal trade-offs for $k$-median clustering in the dynamic setting [HK20, FLNFS21, DHS24, BCG+24, BCF25]. Bhattacharya, Costa, and Farokhnejad [BCF25] achieve state-of-the-art guarantees, obtaining an $O(1)$-approximation with $\tilde{O}(k)$ update time. However, this $\tilde{O}(k)$ is not sublinear time in the worst case (i.e., $k=O(n)$). In contrast, our algorithm always runs in sublinear small polynomial update time, at the expense of incurring a logarithmic factor in the approximation ratio. Previously, no known algorithm can achieve such sublinear update time with any $\operatorname{poly}\log n$ ratio.
We further consider the classic Euclidean bipartite matching problem in the dynamic setting. Although this problem has been recently studied in low dimensions [GKP+25], the update time of their approach exhibits an inherent exponential dependency on the dimension. We use Theorem 1.1 to give the first high-dimensional algorithm for this problem.
Corollary 1.3 (Dynamic Euclidean Bipartite Matching; see Corollary 5.5).
@@
Theorem 1.4 (MPC Tree Embedding; see Theorem 4.3).
@@ 1.2 Technical Contributions
We provide a high-level picture of our technical contributions as follows. Then in Sections 1.2.1 and 1.2.2 we give an overview of proofs, where we talk about the concrete technical challenges and our solution.
@@
To utilize the locality guarantee of Item 1, we employ a metric decomposition called sparse partitions [JLN+05, Fil24]. Specifically, in Euclidean $\mathbb{R}^{d}$, it serves as a $\tau$-BD such that any radius-$r/\Gamma$ metric ball intersects at most $n^{1/\operatorname{poly}(\Gamma)}$ parts. In our applications, we need to use consistent hashing [CJK+22, FJL+25] which is data-oblivious and time/space efficient construction of such sparse partition. This enables us to precisely simulate our new CKR algorithm in the MPC setting. We further show that the embedding undergoes $\tilde{O}(1)$ expected well structured changes under point insertions and deletions. In turn, our new tree embedding results open a door to rich applications. Notably, we devise new dynamic algorithms based on this new tree embedding maintaining optimal solutions with respect to the dynamic embedding tree.
@@ 1.2.1 Proof Overview: Tree Embedding via General Metric Decomposition
In our modified CKR, apart from the input dataset $P\subseteq V$ and scale $w>0$, our algorithm additionally takes a parameter $\Gamma>0$ as input. As mentioned, this parameter balances the accuracy with the locality. Our algorithm start with imposing an $O(w)$-BD/hashing. Then a natural change we make to original CKR, is to replace finding the minimum $\pi$ value in $B^{P}(p,r)$, with that in the intersecting buckets $\tilde{B}(p,r):=\operatorname{buks}^{P}(B(p,r))$, i.e., $\ell_{p}:=\pi_{\min}(\tilde{B}^{P}(p,r))$, where $\pi_{\min}(S):=\min_{x\in S}\pi(x)$. The other related change is that $r$ is picked uniformly at random from $w/\Gamma/4,w/\Gamma/2]$, scaled by $\Gamma$. Finally, there is also a subtle change to defining tree embedding: for each distance scale $2^{m-i}$, we run the modified CKR not only using the scale $w_{i}=2^{m-i}$, but also a new instance of $O(w_{i})$-BD (over $i$) which may not have any relation to other $2^{j}$-BD’s for $j\neq i$, especially that they are not nested/hierarchical. This means we must also define $\tilde{B}^{P}$ as $\tilde{B}^{P}_{i}$ for scale $w_{i}$, instead of using a universal BD. See Algorithm 1 for the full description of this process.
@@
To adapt this FRT’s analysis in our case, a major technical challenge is caused by using $\tilde{B}_{i}(\cdot,\cdot)$ to define the labels for each scale $w_{i}$ (instead of using the metric ball $B(\cdot,\cdot)$). This approximate metric ball $\tilde{B}_{i}(p,r)$ is structurally sophisticated and does not share nice geometric properties as in an exact metric ball, especially considering we are working with an arbitrary bounded decomposition. To see this, given that a point $q$ belongs to $\tilde{B}_{i}(p,r)$, it does not imply a closer point $q^{\prime}$ (to $p$) is also in $\tilde{B}_{i}(p,r)$, and this breaks the probability $\Pr[\mathcal{E}_{j}]\leq 1/j$ as in (1). We emphasize that these challenges are independent to the new parameter $\Gamma$, and the technical challenges remain even if $\Gamma=1$.
@@
To resolve this issue, our main insight is that it is possible to define a set of representative points $\operatorname{rep}(p,r)\subseteq V$ for $\tilde{B}(p,r)$ (and $\operatorname{rep}(p,r)$ may not be a subset of dataset $P$), defined such that the algorithm may be roughly interpreted as running on $\operatorname{rep}(p,r)$, and that it behaves similarly enough to the metric ball $B(p,r)$. This holds for any metric decomposition. In particular, it satisfies the following (see Definition 3.6) and the existence of such representatives is shown in Lemma 3.7, which is our major technical insight.
@@
In this way, we can use $\operatorname{rep}_{i}^{P}(p,r)$ to replace $B^{P}(p,r)$ when computing the label $\ell^{(i)}_{p}$; specifically, the label $\ell^{(i)}_{p}$ can be equivalently found only in $\operatorname{rep}_{i}(p,r)$, instead of the whole $B(p,r)$, i.e., there is $\operatorname{rep}^{*}_{i}(p,r)\in\operatorname{rep}_{i}(p,r)$ such that $\ell^{(i)}_{p}=\pi_{\min}(\operatorname{buks}^{P}_{i}(\operatorname{rep}^{*}_{i}(p,r)))$ (see Definition 3.8). Moreover, we can show an analogy to (1): for some scale $w_{i}$, list $\operatorname{rep}_{i}(p,r)$ in non-decreasing order of distance to $p$ as $p_{j}$’s,
$\forall j,\quad\Pr[\pi_{\min}(\operatorname{buks}_{i}^{P}(p_{j}))=\ell^{(i)}_{p}]\leq H_{m_{j}}-H_{m_{j-1}},$ (2)
where $H_{t}:=\sum_{s=1}^{t}1/s$ is the harmonic sum, and $m_{t}:=|\bigcup_{s\leq t}\operatorname{buks}^{P}(p_{s})|$ is the size of the first $t$ buckets (see (17) in Section 3.2).
In the existence proof of $\operatorname{rep}(\cdot,\cdot)$ (in Lemma 3.7), we actually give an explicit definition of it (used only in the analysis). This definition is technical, and here we discuss the intuition. A natural attempt to define $\operatorname{rep}(\cdot,\cdot)$ such that it satisfies all three of distinct, monotone and ball-preserving properties, is to let $\operatorname{rep}_{i}(r,p):=B(p,r)\cap\operatorname{buks}_{i}(P)$, which is precisely the first half of the “ball-preserving” bullet. The issue is that it does not satisfy the distinctness (bullet 1). Our definition fixes this by only selecting a subset of $B(p,r)\cap\operatorname{buks}_{i}(P)$, roughly the nearest neighbors from $p$ in each bucket of $B(p,r)\cap\operatorname{buks}_{i}(P)$. This “nearest” property ensures the distinctness (bullet 1). Moreover, it is important to select the nearest neighbor and it does not work if e.g., selecting a point in the buckets with smallest ID. The reason is that nearest neighbor has additional geometric properties that ensure bullet 2 and 3 hold.
In fact, we may use a more straightforward statement of (2): list the buckets that intersect $B(p,r)$ in close-to-far order from $p$, and if we denote the $j$-th bucket (in this order) as $S_{j}$, then $\Pr[\ell_{p}^{(i)}=\pi_{\min}(S_{j})]\leq H_{m_{j}}-H_{m_{j}-1}$. However, we still need this explicit identification of representative points, as well as the abstract properties listed in the three bullets, in proving (2), and the other parts of the analysis where the argument is inherently designed to work with points instead of buckets (see e.g., (14), (16)).
@@
Our final analysis deviates from that of FRT, and we indeed need a sum over $i$’s. To avoid the additional $O(\log n)$ factor, we derive a refined upper bound with respect to the scale $w_{i}$, instead of a universal one. The more difficult case is when $w_{i}$ is much larger than $\operatorname{dist}(p,q)$, since if $\ell^{(i)}_{p}\neq\ell^{(i)}_{q}$ happens it incurs a large $w_{i}$ as error, which can be sensitive. For this case, we have an improved bound $O(\operatorname{dist}(p,q))\cdot(H_{\tilde{B}_{i}(p,w_{i})}-H_{\tilde{B}_{i}(p,w_{i}/4)})$ (see Lemma 3.3).
A natural next step is to cancel the terms $H$ out via a telescoping sum when summing over $i$. However, this cancellation requires us to relate $H_{\tilde{B}_{i}}$ with $H_{\tilde{B}_{j}}$ for $j\neq i$. Although the decompositions in different scales may not be related in any way, we can still use the diameter bound and triangle inequality to show that, when the scales differ by a $\operatorname{poly}(\Gamma)$ factor, the smaller-scale approximate ball $\tilde{B}^{P}$ would be a subset of the larger one, regardless of how the decomposition behaves; namely, $\tilde{B}^{P}_{i+s}(p,w_{i}/2^{s})\subseteq\tilde{B}^{P}_{j}(p,r_{i})$ for any $i,j$ and $s=O(\log\Gamma)$ (see Lemma 3.4). This helps us to only incur an additional factor of $O(\Gamma\log\Gamma)$ instead of $O(\log n)$.